Excersise 1- 10 years ago, your faughter bought 4(four),7% semi- annual coupon bond with a 25 year maturity, at the face value of 5000$. How much he will get for these four bonds if he sells them today, knowing that the current interest rate is 6%?


Excercise 2- Last year the company ABC revenues from sales of bicycles were 900.000$, and the average price of the bicycle was 300$. The terms of payment was 45 days net while variable to fixed cost ratio was 80:20. The company has done some financial investments gaining 20% rate of return im stocks trading. This year's the company board members want to increase sales of bicycles and thus suggest relaxing of credit terms from 45 days to 60 days to all buyers projecting the increase in sales for additional 12%.

Calculate and demonstrate your calculation showing if the relaxing kf the credit terms will be acceptable, having in mind the opportunity in investing in profitable stocks with return of 20%.


1
Expert's answer
2021-06-16T09:22:15-0400

1:

The bond price at a given point shows a sum of present values of cash flows expected in the remaining years to maturity. A bond's cash flows include periodic coupon payments and the face value at the maturity date.

In this case, we need to determine the bond selling price today in order to know the total proceeds from selling bonds.

Calculating the number of semi-annual periods in the remaining years to maturity (n):

n=(total years to maturity-years expired) ×Number of semi annual periods per year"=(25-10)×2\\=30"

Calculating the semi-annual coupon payment (C):

C=face value×[annual stated rate ÷number of semi annual periods per year]

"=\$5000×\frac{0.07}{2}\\=\$175"

Calculating the semi-annual discount rate (r):

r=annual market rate ÷number of semi annual periods per year

"=\frac{0.06}{2}\\=0.03 or 3\%"

Calculating the bond price:

"Bond \space price=[C×\frac{1-(1+r)^{-n}}{r} ]+[\frac{face \space value}{(1+r)^{n}}]"

"=[\$175×\frac{1-(1+0.03)^{-30}}{0.03} ]+[\frac{\$5000}{(1+0.03)^{30}}]\\=\$5490.01"

Calculating total proceeds from selling four bonds:

Bond proceeds=bond price × number of bonds sold

"=\$5490.01×4\\=\$21960.04"

Thus, he will get $21,960.04.


2:


The question is related to Management of Receivables. The details are given

Last year's Sales = $900000

Selling Price = $300

Sales in units = $900000 ÷ $300 = 3000 units 

Variable Cost to Fixed Cost ratio = 80:20

Variable Cost = 80

Fixed Cost     20

0.20 Variable Cost = 0.80 Fixed cost 

Fixed Cost = 0.20 Variable Cost 

                 0.80

Fixed Cost = Variable Cost 

               0.40

Variable Cost = 0.40 Fixed Cost. 

Credit period = 45 days.

Current Year 

Sales = 3000 units + 12% × 3000 units   

             = 3360 units 

Credit Period = 60 days 

The Company has an opportunity to invest the amount in some financial investments gaining 20% rate of return im stocks trading. We have the evaluate the credit policy for the current year with regards to last year policy and the investment opportunity. 


Evaluation of Credit Policy

Particulars         Last Year     Current Year 

Sales in units        3000          3360

Selling Price         300           300

Sales in $         900000        1008000

Credit Period        45 days       60 days 

Receivables     900000 × 45   1008000 × 60

                   360           360

Receivables     112500          168000

Increase in Receivables           55500

Desired Return 20% 

Desired Return = 20% × 55500 = $11100

If the investment opportunity provide $11100 as minimum then the Current year policy should be acceptable. 

Note:- For Computational purpose 1 Year taken as 360 days.